OpenStudy (anonymous):

Write the quadratic equation in general form. What is the value of b2 - 4ac?

OpenStudy (bahrom7893):

it's the discriminant..

OpenStudy (anonymous):

what?

OpenStudy (bahrom7893):

what do you mean what value? ax^2 + bx + c = 0 D = b^2 - 4ac

OpenStudy (anonymous):

the general form of the quadratic equation is \[ax ^{2}+bx+c=0\]

OpenStudy (anonymous):

\[b ^{2}-4ac\] is called the Discriminant it value tells us how many and what kind of roots will the given equation have....

OpenStudy (anonymous):

y = (x - 1)2 - 2 y = (x - 1)2 + 2 y = (x + 1)2 - 2 y = (x + 1)2 + 2 theze are the answer choics

OpenStudy (anonymous):

its a dicriminant b^2-4ac wher if \[b ^{2}-4ac\]=0 therefore it has real and equal roots \[b ^{2}-4ac\]<0 it has real and distinct roots \[b ^{2}-4ac\]>0 it has no real roots

OpenStudy (anonymous):

what is your original quadratic equation??? u hv not given that.....

OpenStudy (anonymous):

1 = 2x2 + 7x

OpenStudy (anonymous):

see this equation can be put in the quadratic form as follows: 2x^2 + 7x - 1 = 0 here a = 2, b= 7 and c=-1 so b^2-4ac = (7)^2 - 4(2)(-1) = 49 +8 = 57 so b^2-4ac = 57 Understood????